# Fork, Explore, Commit: OS Primitives for Agentic Exploration

### v1

@@ 2.3. Limitations of Existing Mechanisms
Union filesystems such as UnionFS (Wright et al., 2006) and its successor OverlayFS (Brown, 2016) provide a single writable upper directory with no native commit-to-parent or sibling invalidation, and mounting requires root privileges. Btrfs (Btrfs Project, 2023) and ZFS (Bonwick et al., 2003) support nested subvolumes and clones but are each tied to a single filesystem type and lack native commit-to-parent. Device-mapper snapshots (Linux Kernel Documentation, 2023) require raw block devices with O(depth) read latency. DAXFS (Cong Wang, 2026) provides branching with commit and sibling invalidation but targets memory-backed storage and is not portable.
@@
Feature OverlayFS/UnionFS Btrfs/ZFS DM-Snap DAXFS BranchFS
@@ 5.5. Filesystem-Agnostic Integration
branch() integrates with branching filesystems entirely through three generic ioctls (FS_IOC_BRANCH_CREATE, _COMMIT, _ABORT), following the pattern of existing generic ioctls such as FICLONE and FIEMAP. For FUSE-based filesystems like BranchFS, ioctls are forwarded to the daemon via FUSE_IOCTL; for kernel-native filesystems like DAXFS (Cong Wang, 2026), they are handled directly. Crucially, branch() itself contains no filesystem-specific code: adding a new branching filesystem requires only implementing the three ioctls, with no changes to branch() or the VFS layer.
@@ 8. Discussion and Future Work
Current agents (Claude Code, SWE-agent, OpenHands) primarily checkpoint via files rather than process memory, suggesting that BR_FS-only mode covers most current use cases. The BR_MEMORY flag provides future capability for agents with significant in-memory state (e.g., persistent Python interpreters or embedding caches). Integration requires minimal changes: the orchestrator creates branches and commits the winner, while agents run unmodified inside their branch’s mount namespace.
@@ 9. Conclusion
Agentic exploration requires isolating both filesystem and process state across parallel execution paths, with atomic commit and rollback semantics. We introduce the branch context abstraction to capture these requirements and present two complementary implementations in Linux. BranchFS is a FUSE-based filesystem that provides lightweight copy-on-write branching with O(1) creation and atomic commit-to-parent, all without root privileges. branch() is a proposed Linux syscall for kernel-enforced process coordination with reliable termination and first-commit-wins semantics, integrating with branching filesystems through generic FS_IOC_BRANCH_* ioctls that keep the syscall filesystem-agnostic. Preliminary evaluation shows O(1) branch creation and modification-proportional commit costs.

### v2

@@ 2.3. Limitations of Existing Mechanisms
Union filesystems such as UnionFS (Wright et al., 2006) and its successor OverlayFS (Brown, 2016) provide a single writable upper directory with no native commit-to-parent or sibling invalidation, and mounting requires root privileges. Btrfs (Btrfs Project, 2023) and ZFS (Bonwick et al., 2003) support nested subvolumes and clones but are each tied to a single filesystem type and lack native commit-to-parent. Device-mapper snapshots (Linux Kernel Documentation, 2023) require raw block devices with O(depth) read latency.
@@
Feature OverlayFS/UnionFS Btrfs/ZFS DM-Snap BranchFS
@@ 4.4. Mount-Based Isolation and Interface
4.5. Agent Integration Library
To bridge BranchFS with agent frameworks, we provide BranchContext (Wang and Zheng, 2026), a Python library that wraps BranchFS’s branching primitives into high-level exploration patterns. At its core, a Workspace context manager mounts BranchFS and exposes Branch objects with automatic commit-on-success and abort-on-failure semantics. On top of this, the library provides seven composable patterns that directly realize the agent strategies described in Section 2: Speculate races $N$ candidates and commits the first success; BestOfN runs $N$ candidates and commits the highest-scoring one; Reflexion retries sequentially, feeding failure feedback into the next attempt; TreeOfThoughts explores hierarchical strategy trees with nested branches; BeamSearch keeps the top-$K$ branches alive at each depth level; Tournament performs pairwise elimination via a judge function; and Cascaded starts with one attempt and adaptively fans out on failure. Each pattern manages branch lifecycle (creation, commit, abort) and process isolation (via fork(2) with optional timeout and resource limits) internally, so agent code only supplies the per-branch task logic.
@@ 5.5. Filesystem-Agnostic Integration
branch() integrates with branching filesystems entirely through three generic ioctls (FS_IOC_BRANCH_CREATE, _COMMIT, _ABORT), following the pattern of existing generic ioctls such as FICLONE and FIEMAP. For FUSE-based filesystems like BranchFS, ioctls are forwarded to the daemon via FUSE_IOCTL; for kernel-native filesystems, they are handled directly. Crucially, branch() itself contains no filesystem-specific code: adding a new branching filesystem requires only implementing the three ioctls, with no changes to branch() or the VFS layer.
@@ 8. Discussion and Future Work
Current agents (Claude Code, SWE-agent, OpenHands) primarily checkpoint via files rather than process memory, suggesting that BR_FS-only mode covers most current use cases. The BR_MEMORY flag provides future capability for agents with significant in-memory state (e.g., persistent Python interpreters or embedding caches). The BranchContext library (Section 4.5) already provides this integration for Python-based agents, wrapping BranchFS into ready-to-use exploration patterns so that agent code only supplies per-branch task logic. Agents run unmodified inside their branch’s mount namespace; the library manages branch lifecycle and process isolation.
@@ 9. Conclusion
Agentic exploration requires isolating both filesystem and process state across parallel execution paths, with atomic commit and rollback semantics. We introduce the branch context abstraction to capture these requirements and present two complementary implementations in Linux. BranchFS is a FUSE-based filesystem that provides lightweight copy-on-write branching with O(1) creation and atomic commit-to-parent, all without root privileges. On top of BranchFS, the BranchContext Python library provides seven composable exploration patterns (parallel speculation, best-of-N, reflexion, tree-of-thoughts, and others) for direct integration with agent frameworks. branch() is a proposed Linux syscall for kernel-enforced process coordination with reliable termination and first-commit-wins semantics, integrating with branching filesystems through generic FS_IOC_BRANCH_* ioctls that keep the syscall filesystem-agnostic. Preliminary evaluation shows O(1) branch creation and modification-proportional commit costs.